OpenStudy (anonymous):

A drawer contains five socks, three red and two blue. If you pick a sock randomly, replace it, and pick another, what is the probability that both socks are red? If you pick two socks randomly without replacing the first, what is the probability that both are red?

OpenStudy (anonymous):

In the first case, it is two independent events... very nice.

OpenStudy (anonymous):

So to get one sock it is \[ \frac{3}{5} \]And two get two socks it is \[ \left(\frac35\right)^2 = \frac9{25} \]

OpenStudy (anonymous):

In the second case, after you pull a red sock, probability changes...

OpenStudy (anonymous):

So for the first sock it is \[ \frac35 \]But now you have \(3-1=2\) red socks and \(5-1=4\) total socks So you get \[ \frac24 \]Probability for second try if you know the first try was red.

OpenStudy (anonymous):

Fortunately this 'if' probability is independent from the first sock drawing, which makes the events independent again! So the combined probability is \[ \frac35 \times \frac24=\frac6{20}=\frac3{10} \]

OpenStudy (anonymous):

Thank you!